OpenStudy (anonymous):

The per-copy cost of printing, marketing, and shipping x million copies of Danielle Tinne's new novel is given by the formula C(x)=753-225x+3x^3 Find the number of copies that lead to the lower cost per copy, and find that minimum cost.

OpenStudy (anonymous):

@AccessDenied @Whiterabbit541 @danielbarriosr1 @francesco.alemanno PLEASE HELP ME

OpenStudy (rational):

straightforward problem, find the derivative and set it equal to 0

OpenStudy (anonymous):

how do i do that. Can you show me step by step

OpenStudy (rational):

know how to take derivative of a function ?

OpenStudy (anonymous):

can i switch the numbers around like 3x^3-225x+753

OpenStudy (rational):

definitely :)

OpenStudy (anonymous):

Okay so C(x)=\[3x^3-225x+753\]

OpenStudy (anonymous):

the derivative is 9x^2-225

OpenStudy (anonymous):

@rational

OpenStudy (rational):

excellent ! set it equal to 0 and solve x

OpenStudy (anonymous):

9x^2-225=0

OpenStudy (anonymous):

wait how do i solve for x if theres an exponent

OpenStudy (rational):

first cancel out 9

OpenStudy (rational):

9x^2 - 225 = 0 9(x^2 - 25) = 0 x^2 - 25 = 0

OpenStudy (rational):

x^2 = 25 x = ?

OpenStudy (anonymous):

isnt it suppose to be 225 where did you got 25?

OpenStudy (anonymous):

@rational

OpenStudy (rational):

I have factored out 9 from 225 : 225 = 9*25

OpenStudy (anonymous):

and what happens to the 9?? you just take it out?

OpenStudy (rational):

Alright, forget about my earlier work

OpenStudy (rational):

Here is another way to simplify : 9x^2 - 225 = 0 9x^2 = 225 x^2 = 225/9

OpenStudy (rational):

whats the value of 225/9 ?

OpenStudy (rational):

whats the value of 225/9 ?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

now we need to find the minimum cost

OpenStudy (rational):

yes minimum value occurs at x = 5

OpenStudy (anonymous):

But whats the minumum cost? thats the question

OpenStudy (rational):

To get the minimum cost, evaluate the given function at x = 5

OpenStudy (rational):

C(x)=753-225x+3x^3 C(5) = ?

OpenStudy (anonymous):

from the original function or the derivative

OpenStudy (rational):

from the original function only. the job of derivative is over. it just tells u where the minimum/maximum value OCCURS.

OpenStudy (anonymous):

minimum cost is 3

OpenStudy (rational):

to get the minimum/maximum value, u need to evaluate the original given function

OpenStudy (rational):

3 is correct !

OpenStudy (rational):

so, for minimum cost she needs to make 5 million copies and the minimum cost is 3 whatever unit it is